OpenStudy (anonymous):

URGENT! Write an equation in point-slope form for the line through the given point with the given slope. (10,-9); m=-2

OpenStudy (mathstudent55):

You need an equation in this form: \(\large y = mx + b\)

OpenStudy (anonymous):

okay so do you know y=mx=b

OpenStudy (anonymous):

y=mx+b

OpenStudy (mathstudent55):

Let m = the given slope, so m = -2. We now have: \(\large y = -2x + b\)

OpenStudy (anonymous):

sorry ;/

Nnesha (nnesha):

*point slope* \[\huge\rm y-y_1=m(x-x_1)\]

imqwerty (imqwerty):

i jst came here to smile :)

OpenStudy (mathstudent55):

Now put the x and y of your point in the x and y of our equation and solve for b. \(\large y = -2x + b\) \(\large -9 = -2(10) + b\) \(\large -9 = -20 + b\) \(\large 11 = b\) Now that we know b = 11, substitute 11 for b. \(\large y = -2x + b\) \(\large y = -2 x + 11\)

OpenStudy (mathstudent55):

@Nnesha is correct above. They ask for point-slope, not slope-intercept. \(\large y - y_1 = m(x - x_1) \) \(\large y - (-9) = -2(x - 10)\)